# Robot Localization aided by Quantum Algorithms

### v1

@@ 4.3 More complex grid, with a pattern match with two qubits
To thoroughly assess the performance and robustness of the proposed quantum algorithm for robot localization, we designed a setup that pushes the algorithm to its limits. In this case, a 4x4 grid will be used, with a single qubit per cell (the calculated costmap in each cell has been configured to use a single qubit). Once again, the Pygame based interface has been used to configure such grid, and to define the robot’s position in a certain cell in the grid.
The robot has been positioned in an arbitrary location, as seen in Figure 17.
Map qubits: $\begin{bmatrix}1&0&0&0\\ 1&0&1&0\\ 1&1&0&0\\ 0&0&0&0\end{bmatrix}\$
Horizontal search qubits: $\begin{bmatrix}1,0\end{bmatrix}\\ \$
Vertical search qubits: $\begin{bmatrix}1,0\end{bmatrix}\$
Figure 17: : Simple setup with a 4x4 grid
@@
The proposed double Grover search with a pattern $[1,0]$ in a 4x4, has to deal with a search space of $33x$ 3 (as the length of the pattern to be searched requires matching two cells). In order to evaluate the quantum circuit, it’s necessary to use:
  • 4 qubits for the search space (to encode positions 0 to 9)
@@
  • The Oracle has to deal with 9 possible positions (3x3), with an expectation of a single correct solution (N=9, M=1)
  • The suggested number of repetitions is: $R=\pi/4\sqrt{N/M}=2$
  • The total number of required qubits is 27
Once again, the quantum circuit code has been tested: locally in a computer (using qasm), using the IonQ cloud, the Bluequbit cloud, and in a real quantum computer (from IBM). It was not possible to simulate the circuit on QuantumInspire as the total number of required qbits (27) exceeded the capacity of the simulated environment (26qubits). Results are shown in Fig. 18.
(a) QasmSimulator
(b) IonQ
(c) Bluequbit
(d) Real IBM Quantum computer
@@
Figure 18: : Histograms of the same quantum circuit obtained from simulated and real quantum computers
@@
In simulations, Grover’s algorithm clearly identified the correct solution in the simulated circuits: position 4 (0010) of all possible locations, which means row 2 and column 2 (count starts from zero). But although in all the simulated tests, the correct solution’s probability was clearly amplified, this did not happen in the real quantum computer.
@@
Using again code from (Aseginolaza et al. (2024)), for this 4x4 grid we get:
  • Error due to noisy gates = 100%
  • Error due to estability $>99.99\%$
  • Measurement errors $>67\%$
  • Total estimated error probability = 100%
@@
Currently, real quantum hardware lacks the robustness and scalability needed for practical applications, limiting its use in real situations.
@@
Nevertheless, researchers are actively working to overcome these challenges, developing new quantum error correction techniques and improving qubit coherence times. While noise mitigation techniques are a crucial aspect of reliable quantum computing, a detailed exploration of these methods is beyond the scope of this paper. The application of noise mitigation techniques, such as dynamic decoupling, noise reduction, and quantum error mitigation, requires a comprehensive understanding of quantum error correction codes, quantum control techniques, and classical algorithms.
@@ 5 Conclusions and Future Work
The experimental results obtained from the proposed quantum algorithm for robot localization demonstrate its feasibility and effectiveness in solving the problem. The approach successfully localizes the robot in a 2D map, confirming the theoretical validity of the quantum-based method. However, the performance of the algorithm on real quantum computers is hindered by the current technological limitations. As the size of the map increases, the algorithm’s accuracy degrades due to the difficulties in maintaining the fragile quantum state for extended periods. This is a common challenge faced by many quantum algorithms, and our results highlight the need for advancements in quantum hardware to fully exploit the potential of quantum computing.
Furthermore, our simulations on classical computers reveal another limitation: the number of qubits required to simulate the algorithm grows exponentially with the size of the map. Currently, we are unable to simulate the algorithm for maps that require more than 40 qubits, which restricts the scalability of our approach. This emphasizes the importance of developing more efficient quantum algorithms or simulation methods that can overcome these limitations.
Despite the promising results, there are several research lines for future work to improve the performance and applicability of the proposed quantum algorithm.
  • Non-exact match quantum algorithm : One potential direction is to explore non-exact match quantum algorithms, such as those based on Hamming distance search. This could allow for more robust and flexible localization, especially in the presence of noisy or incomplete data.
@@
  • Quantum error correction and noise mitigation : to overcome the challenges posed by current quantum hardware, researching and implementing quantum error correction techniques or noise mitigation strategies could help maintain the coherence of the quantum state for longer periods.
  • Hybrid quantum-classical approaches : exploring hybrid approaches that combine the strengths of quantum and classical computing could provide a more practical and efficient solution for robot localization. This might involve using quantum computing for specific sub-tasks or exploiting classical machine learning techniques to supplement the quantum algorithm.

### v2

@@ 4.3 Grid with a pattern match with two qubits
To thoroughly assess the performance and robustness of the proposed quantum algorithm for robot localization, we designed a setup that explores the use of patterns longer than 1 qubit. In this case, a 4x5 grid will be used, with horizontal and vertical pattern searches that use two qubits. Once again, the Pygame based interface has been used to configure such a grid and to define the robot’s position in a certain cell in the grid.
An obstacle has been placed in an arbitrary location, as seen in Figure 17.
Map qubits: $\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&1&0\\ 0&0&1&1&0\end{bmatrix}\$
Horizontal search qubits: $\begin{bmatrix}1,1\end{bmatrix}\\ \$
Vertical search qubits: $\begin{bmatrix}1,1\end{bmatrix}\$
Figure 17: Setup with a 4x5 grid
@@
The proposed double Grover search with a pattern $[1,1]$ in a 4x5, has to deal with a search space of $43x$ 4 (as the length of the pattern to be searched requires two matching cells). In order to evaluate the quantum circuit, it is necessary to use:
  • 4 qubits for the search space (to encode positions 0 to 12)
@@
  • The Oracle has to deal with 12 possible positions (4x3), with an expectation of a single correct solution (N=12, M=1)
  • The suggested number of repetitions is: $R=\pi/4\sqrt{N/M}=3$
  • The total number of required qubits is 31
  • Circuit depth is 292
As this setup required 31 qubits, the system could not be tested locally on a computer. The only options available are to use the Bluequbit cloud, or run the algorithm in a real quantum computer.
The results are shown in Fig. 18.
(a) Bluequbit simulation
(b) Results of running the algorithm in a real ’ibm kyiv’ quantum unit
@@
Figure 18: Histograms of the same quantum circuit obtained from simulated and real quantum computers
@@
In simulations, Grover’s algorithm clearly identified the correct solution in the simulated circuits: position 10 (0101) of all possible locations, which means row 2 and column 2 of the search space (count starts from zero). But although in all the simulated tests the correct solution’s probability was clearly amplified, this did not happen in the real quantum computer.
An approximation for the probability of having an erroneous result using the ’ibm kyiv’ quantum computer is:
$\begin{split}P(\text{at least 1 gate fails})&=1-(1-3.81e-3)^{292}\\ &\approx 67\%\end{split}$ (6)
@@
Using again the code from ((Aseginolaza et al., 2024)), for this $4\times 5$ grid, we get:
  • Error due to noisy gates $>99\%$
  • Measurement errors $>50\%$
@@
It seems clear that current real quantum hardware lacks the robustness and scalability needed for practical applications, limiting its use in real situations.
@@
Nevertheless, researchers are actively working to overcome these challenges, developing new quantum error correction techniques and improving qubit coherence times. Although noise mitigation techniques are a crucial aspect of reliable quantum computing, a detailed exploration of these methods is beyond the scope of this paper. The application of noise mitigation techniques, such as dynamic decoupling, noise reduction, and quantum error mitigation, requires a comprehensive understanding of quantum error correction codes, quantum control techniques, and classical algorithms.
5 Scalability
As mentioned, the quantum localization algorithm introduced in this paper proposes a novel method to reduce the computational complexity of searching for a robot’s position in an $n\times n$ grid. As the complexity of the localization problem increases with the size of the grid (denoted as N), the proposed method takes advantage of an approach inspired by Grover’s algorithm to reduce this complexity to ${O(\sqrt{N})}$.
Although the algorithm is theoretically unrestricted, its practical implementation remains limited by current engineering constraints, as demonstrated in previous experimental setups. Presently, quantum computers face capacity restrictions due to noise and the inherent difficulty of maintaining quantum states over extended periods. These disturbances can introduce errors in quantum computations, posing significant challenges in the execution of reliable operations.
However, it should be noted that this situation is expected to change in the future. Quantum gate error rates are expected to be significantly reduced in the coming years due to advances in error correction techniques, improved qubit coherence times, and enhanced fabrication processes, as shown in the trend presented in Figure 19
Figure 19: Recent progress in quantum gate error rates. Source: https://nap.nationalacademies.org/read/25196/chapter/9
To empirically assess the current limits of running the proposed algorithm on real quantum computers, a series of test sets have been prepared and executed in both simulated environments (locally and through third parties) and real quantum hardware provided by IBM via the IBM Quantum Platform.
These tests are presented as a simple problem of finding the position of an obstacle (marked with a "1") in a certain grid. These grids scale from $1\times 2$ (2 elements) to $6\times 6$ (64 elements). The results are shown in Figure 21.
As an additional comment, it should be noted that empirical testing has also shown that the complexity of the elements within the grid does not affect the performance of the search algorithm. In other words, the difficulty of finding the robot location remains the same, regardless of whether the grid contains a single distinguishing feature or multiple features. This means that the grids shown in Figure 20 require essentially the same amount of computation.
Figure 20: Grids of the same size with different amount of features require the same amount of computation
Figure 21: Scalability test results
Note that for the tests shown in Figure 21, an error probability of 3.81e-3 has been considered (in line with what can be achieved in current modern quantum computers), and that the error has been calculated using the methodology proposed by (Aseginolaza et al., 2024).
6 Quantum advantage
Quantum advantage is the situation at which a quantum computer can solve a problem faster or more efficiently than the best classical computers available.
This does not necessarily mean that the quantum computer is nowadays useful for practical applications in general: it just states that it outperforms classical systems on a specific task (in other words: although these breakthroughs mark important milestones, they do not yet necessarily translate into general-purpose, practical applications).
Quantum advantage is typically demonstrated in highly specialized problems, such as the localization problem presented in this paper.
Today in robotics, probabilistic methods are commonly employed to address the localization problem, as they can effectively combine the strengths of various localization techniques, mitigate their weaknesses, and handle uncertainty and noisy measurements, as noted in (Bongard05). That is why in this paper we compare the efficiency of the proposed quantum-aided localization algorithm to a traditional probabilistic robot localization method in a grid-based scenario.
Probabilistic methods include different approaches, such as Bayesian filtering, such as Kalman filters, Extended Kalman filters, Unscented Kalman filters, Monte Carlo localization (MCL), etc. The choice of probabilistic method depends on the specific application, sensor suite, and performance requirements. Some methods, like MCL, are more computationally expensive but provide more accurate estimates, while others, like Kalman filters, are more efficient but may be less accurate.
As one of such probabilistic techniques, in this comparison Monte Carlo Localization (MCL) is used. MCL uses random sampling to represent the robot state as a set of particles. Each particle is weighted based on its likelihood, and the weights are updated over time to reflect the changing environment and sensor measurements, so for localization, the typical workflow would be:
  • Initialize: Start with a uniform belief over all cells (if no prior knowledge).
  • Move: If the robot is in motion, apply the motion model to predict the new state.
  • Sense: Compare sensor readings with expected values and update probabilities.
  • Normalization: Normalization of measured values.
  • Resample: remove unlikely positions and focus on probable ones.
If we take, for example, the grid shown in Figure 17 (a 4x4 grid), we have an uncertainty area of 9 cells (since the setup is using two qubits). As MCL particles are initially spread uniformly and randomly throughout the space state, it is expected that (in our case) there are approximately 9 particles that have to: be initialized, sense their surroundings, and then resampled. As the grid grows, more particles are needed to cover the map, so for a map with N cells, the complexity of localization increases with $O(N)$.
Based on the initial code proposal of (RoboND-MCL-Lab), we simulate the configuration (described in Figure 17). Figure 22 shows the evolution of the MCL algorithm to estimate robot location if starting from an unknown location.
(a) Step 0
(b) Step 1
(c) Step 2
(d) Step 3
Figure 22: Iterations of MCL in the setup described in 17. Black squares represent obstacles, green dots the particles (in light green the resampled particles), and the blue circle represents the robot
Figure 22 shows how the particle filter iteratively approaches to the solution, starting from a set of randomly distributed particles.
In our proposed approach, since Grover’s algorithm requires repeating certain steps according to the size of the state map $(N)$ following $Repeat=\pi/4\sqrt{N}$, complexity increases with the square root of N.
This means that for large values of N, O(sqrt(N)) grows much slower than O(N). This means that the second method will be significantly faster for large inputs.
To illustrate this, consider an example:
  • If the size of the map is N = 100, then $\sqrt{N}\approx 10$ for the quantum-aided approach, and N = 100 for classical methods.
  • If N = 1,000,000, then $\sqrt{N}\approx 1,000$, and N = 1,000,000.
As N grows, the relative difference between O(N) and O(sqrt(N)) becomes more pronounced, and O(N) grows much faster. This implies that for larger maps the proposed quantum-aided approach requires much fewer calculations (obviously, assuming there exists a real quantum device that can compute such calculations).
@@ 7 Conclusions and Future Work
The experimental results obtained from the proposed quantum algorithm for robot localization demonstrate its feasibility and effectiveness in solving the problem. The approach successfully localizes the robot on a 2D map, confirming the theoretical validity of the quantum-based method. However, the performance of the algorithm on real quantum computers is hindered by current technological limitations. As the size of the map increases, the accuracy of the algorithm degrades as a result of the difficulties in maintaining the fragile quantum state for extended periods. This is a common challenge faced by many quantum algorithms, and our results highlight the need for advances in quantum hardware to fully exploit the potential of quantum computing.
Furthermore, our simulations on classical computers reveal another limitation: the number of qubits required to simulate the algorithm grows exponentially with the size of the map. Currently, we are unable to simulate the algorithm for maps that require more than 40 qubits, which restricts the scalability of our approach. The limitations of current quantum systems underscore the need for developing more efficient quantum algorithms and simulation methods. These advancements will help bridge the gap until true quantum computers can effectively utilize a larger number of qubits.
Fortunately, recent advances in the design and manufacturing of quantum computers indicate a promising trajectory, with many of these current shortcomings expected to be addressed within the next decade.
Additionally, coherence time (the amount of time that a qubit can maintain its quantum state, or coherence, without being disrupted or affected by its environment) is also being increased following a similar trend. This means that longer (or more complex) calculations can be executed without errors in a certain algorithm.
Figure 23: Available qubits per chip in some of the main quantum computer providers (left) and evolution of coherence times (right). Sources: https://www.netmeister.org/blog/pqc-2024-01.html and Eindhoven University
However, despite the promising results, there are several research lines for future work to improve the performance and applicability of the quantum algorithm proposed in this paper. Work on this paper has been focused in the design and validation of the algorithm, but other aspects can be developed in the future, such as:
  • Non-exact match quantum algorithm: One potential direction is to explore non-exact match quantum algorithms, such as those based on the Hamming distance search. This could allow for more robust and flexible localization, especially in the presence of noisy or incomplete data.
@@
  • Quantum error correction and noise mitigation: To overcome the challenges posed by current quantum hardware, research and implementation of quantum error correction techniques or noise mitigation strategies could help maintain the coherence of the quantum state for longer periods.
  • Hybrid quantum-classical approaches: Exploring hybrid approaches that combine the strengths of quantum and classical computing could provide a more practical and efficient solution for robot localization. This might involve using quantum computing for specific subtasks or exploiting classical machine learning techniques to supplement the quantum algorithm.
@@
During the preparation of this work, the author used Llama3.1 generative AI in order to revise the spelling, grammar errors and for rephrasing sentences for clarity. After using this tool/service, the author reviewed and edited the content as needed and assumes full responsibility for the content of the publication.
This research was funded by the ELKARTEK Research Program of the Basque Government, project #KK-2024/00024. The APC was funded by Fundación TECNALIA Research and Innovation.
The authors declare no conflict of interest. The funders had no role in the design of the study; in the collection, analyses, or interpretation of data; in the writing of the manuscript; or in the decision to publish the results.